# customstyle linecolor=gray!100, linewidth=3pt, innerleftmargin=3pt, topline=false, rightline=false, bottomline=false, leftline=true, innerrightmargin=3pt, innertopmargin=3pt, innerbottommargin=3pt, backgroundcolor=gray!15

### v1

@@ 3. Methodology
We propose ColaUntangle, a collaborative framework that leverages LLM-driven agents to untangle commits by identifying both explicit and implicit dependencies among code changes. As shown in Figure 2, ColaUntangle first extracts structured code change information, including commit diffs and contexts derived from $\delta$-PDG. It then employs a multi-agent consultation process, where agents iteratively analyze dependencies and synthesize untangling decisions until reaching consensus.
@@
Figure 3. Code Change Example with Explicit and Implicit Contexts Derived from $\delta$-PDG
@@ 3.1. Structured Code Change Information
To enrich the contextual information available for dependency analysis, we first extract explicit contexts and implicit contexts from the multi-version program dependency graph ($\delta$-PDG).
@@
A program dependency graph (PDG) (ferrante1987PDG, ) is a directed graph that consists of a set of nodes N and a set of edges E, where each node n $\in$ N represents a program statement or a conditional expression; and each edge e $\in$ E represents the data flow or control flow among the statements. However, PDG can only describe the dependency relationships between statements of individual program. In order to simultaneously reflect the statement dependencies before and after program modification, we also need to introduce $\delta$-PDG. The multi-version program dependency graph ($\delta$-$PDG^{i,j}$) (parțachi2020flexeme, ) is the disjoint union of all nodes and edges across all versions in the PDGs at versions i and j.
For most modification operations, the changed parts account for only a small proportion of the entire program. To reduce unnecessary information, we introduce the concept of explicit contexts. The explicit contexts is a compressed $\delta$-PDG graph obtained by retaining only the modified nodes and their relationships in the $\delta$-PDG. Although the explicit contexts loses some context information, it still retains the data dependencies and control dependencies between the changed nodes. Consequently, the implicit contexts retains only the modified nodes in the $\delta$-PDG and their neighbor nodes in one hop. Compared with the explicit contexts, the implicit contexts retains more surrounding code information.
@@
Figure 3 illustrates a code change scenario involving two statement replacements (at lines 2 and 6). Figure 3 (a) shows the $\delta$-PDG of this code snippet, where red nodes represent the statements before the change, green nodes represent the modified statements, and gray nodes denote the unchanged statements. Figure 3 (b) presents the explicit contexts, which are obtained by compressing the $\delta$-PDG in Figure 3 (a). Specifically, we retain the changed nodes and their direct dependency edges; additionally, if two changed nodes lack a direct dependency but are connected via a dependency path (channel), we add a compressed edge between them. For example, nodes 2 and 6 have no direct dependency but are connected via a path through nodes 4 and 5, resulting in a compressed edge between them. This compressed edge is labeled with the variables involved in the data flow along the original dependency path. Figure 3 (c) shows the implicit contexts, which retain both the changed nodes and their immediate neighbors (within one-hop dependencies). In this example, node 9 is excluded because it does not have a direct dependency with any of the changed nodes.
@@ 3.2.1. Agent Building
Figure 4. Worker Agent
Figure 5. Reviewer Agent
@@
ColaUntangle builds three types of agents. The first two are worker agents, each specializing in untangling commits based on either explicit or implicit dependencies between code changes, as illustrated in Figure 4. The third is a reviewer agent, which specializes in reviewing and synthesizing the results from the worker agents, as shown in Figure 5.
@@
Worker Agent. Based on the explicit and implicit dependency rules summarized in Section 2.2, we define two types of worker agents, as illustrated in Figure 4. The worker agents specialize in untangling commits by analyzing explicit and implicit dependencies between code changes, respectively. We refer to the explicit worker agent as EA and the implicit worker agent as IA. We define the roles of these agents using the following role definition prompt: “You are an expert in untangling tangled commits, specializing in analyzing explicit (or implicit) dependencies between code changes and comprehending their underlying semantic relationships”. We also provide each agent with the corresponding explicit or implicit dependency rules (concluded in Section 2.2.1) to guide their analysis. Each worker agent has two main functions. The first function is “untangle”. EA untangles tangled commits based on the inputs of commit diffs and explicit contexts, while IA performs untangling based on commit diffs and implicit contexts. Both agents are required to generate explanations for their decisions to facilitate understanding of their underlying criteria. The second function is “validate”. In this function, each worker agent compares the synthesized result generated by the reviewer agent (described in the next section) with its own untangling result and provides its opinion on the synthesized result. This includes indicating agreement or disagreement and, if disagreeing, providing a revised untangling result along with an explanation.
@@
Reviewer Agent. The untangling results produced by the two worker agents may conflict due to their different expertise and the presence of diverse explicit and implicit dependencies within code changes. To resolve these conflicts, we define a reviewer agent, referred to as RA, which specializes in reviewing and synthesizing the untangling results from the worker agents, as illustrated in Figure 5. We define the reviewer agent’s role using the following role definition prompt: “You are an expert reviewer specializing in reviewing and synthesizing untangling results from worker agents. You have deep expertise in analyzing both explicit and implicit dependencies between code changes”. To support its role, we provide RA with both explicit and implicit dependency rules to guide its synthesis process. RA has three main functions: synthesize, revise and stop. In the synthesize function, RA synthesizes the results from EA and IA based on the structured code change information (commit diffs, explicit contexts, and implicit contexts). In the revise function, RA revises its synthesized result based on the new opinions generated by the two worker agents during validation. Finally, in the stop function, when the collaborative consultation process reaches its maximum iteration limit, RA produces the final untangling result. For each function, RA is required to provide clear explanations for its decisions.
@@
1:Experts $EA$, $IA$, $RA$; initial synthesized result $r_{RA}$; maximum attempts $t$; prompt template $Prompt$
@@
3:// Initialize variables
4:$consensus\leftarrow False$, $round\leftarrow 0$
5:// Iterative consultation
6:while $consensus=False$ and $round<t$ do
7:    $round\leftarrow round+1$
8:    $consensus\leftarrow True$
9:    $EA\_opinion\leftarrow EA.validate(Prompt(r_{RA}))$
10:    $IA\_opinion\leftarrow IA.validate(Prompt(r_{RA}))$
11:    if $EA\_opinion.agree=False$ or $IA\_opinion.agree=False$ then
12:        $consensus\leftarrow False$
13:        if $round=t$ then
14:           $r_{RA}\leftarrow RA.stop(EA\_opinion,IA\_opinion)$
15:        else
16:           $r_{RA}\leftarrow RA.revise(EA\_opinion,IA\_opinion)$
17:        end if
18:    end if
19:end while
20:$r_{f}\leftarrow r_{RA}$
21:return $r_{f}$
3.2.2. Utangling Workflow
In the first phase, generation of initial reviews, EA and IA are instructed to generate their preliminary untangling results $r_{EA}$ and $r_{IA}$ using their Untangle function, based on the commit diffs ($Info_{diffs}$) and the explicit ($Info_{EC}$) or implicit ($Info_{IC}$) contexts:
$r_{EA}=EA.Untangle(Prompt(Info_{diffs},Info_{EC})$
$r_{IA}=IA.Untangle(Prompt(Info_{diffs},Info_{IC})$
$r_{RA}=RA.Synthesize(Prompt(r_{EA},r_{IA},Info_{diffs},Info_{EC},Info_{IC})$
@@ 3.2.2. Utangling Workflow
$r_{RA}^{j}=RA.revise(Prompt(EA\_opinion^{j},EA\_opinion^{j})$
@@
$r_{RA}^{j}=RA.stop(Prompt(EA\_opinion^{j},EA\_opinion^{j})$
@@ 4.1. Datasets
We use two datasets that have been widely used in previous work: 1) A C# dataset used in (chen2022attributedgraph, ; herzig2016impact, ; parțachi2020flexeme, ; li2022utango, ; fan2024HD-GNN, ), which contains 1,612 tangled commits (each of which has $\leq$ 3 concerns) across 9 GitHub projects. 2) A Java dataset used in (li2022utango, ; fan2024HD-GNN, ). The Java dataset contains over 14k tangled commits in 10 GitHub repositories, where the number of concerns in a commit ranges from 2-32. Both datasets were constructed by first selecting atomic commits based on criteria such as temporal proximity, namespace similarity, and file coupling, and then artificially tangling them via git cherry-picking to simulate realistic tangled commit scenarios.
@@ 4.2. Baselines
  • Barnett et al. (barnett2015Heuristic-rule, ): A heuristic rule-based approach considering def-use, use-use, and same-enclosing-method relations among code changes.
  • Herzig et al. (herzig2016impact, ): A heuristic rule-based approach combine various Confidence Voters and build triangle partition matrix to untangle the commits.
  • Flexeme (parțachi2020flexeme, ): A graph clustering-based approach builds a $\delta$-NFG (proposed in Flexeme) of commits and then applies agglomerative clustering on this $\delta$-NFG to untangle the commits.
  • $\delta$-PDG+CV (parțachi2020flexeme, ): A variant of Flexeme by applying Herzig et al.’s confidence voters directly to $\delta$-PDG (proposed in Flexeme).
  • UTango (li2022utango, ): A Graph clustering-based approach builds a $\delta$-PDG from commits and then applies GNN and agglomerative clustering on this $\delta$-PDG to untangle the commits.
  • HD-GNN (fan2024HD-GNN, ): A Graph clustering-based approach builds a entity reference graph from commits and then applies GNN to aggregate the embeddings of nodes and edges in the graph and classify the relation between the code changes.
The baselines used for Java dataset are: Barnett et al., Herzig et al., UTango, and HD-GNN mentioned above, as all these approaches support analysis of Java program files. SmartCommit is a graph clustering-based interactive method that builds a Diff Hunk Graph and applies a graph partitioning algorithm to untangle commits. We also include SmartCommit and its baselines, Base-1 (a rule-based approach that groups all changes into a single concern) and Base-2 (which groups changes by file), in our comparison.
LLM-based Baselines. We incorporate baselines utilizing single LLM for both C# and Java datasets. We employ two different prompt strategies: zero-shot and zero-shot CoT. Zero-shot refers to prompting LLM to perform a task without providing examples, zero-shot CoT directly incorporates the prompt Let’s think step by step after a question to facilitate inference (kojima2022zeroshot, ).

### v2

@@ 3. Methodology
We propose ColaUntangle, a collaborative framework that leverages LLM-driven agents to untangle commits by identifying both explicit and implicit dependencies among code changes. As shown in Figure 2, ColaUntangle first extracts structured code change information, including commit diffs and contexts derived from source code. It then employs a multi-agent consultation process, where agents iteratively analyze dependencies and synthesize untangling decisions until reaching consensus.
@@ 3.1. Structured Code Change Information
To enrich the contextual information available for dependency analysis, we first extract explicit contexts and implicit contexts from the program dependency graphs (PDGs) from the source code. Our method centers on building a consolidated graph structure sensitive to code changes, followed by targeted reduction and compression.
@@
A program dependency graph (PDG) (ferrante1987PDG) is a directed graph that consists of a set of nodes N and a set of edges E, where each node n $\in$ N represents a program statement or a conditional expression; and each edge e $\in$ E represents the data flow or control flow among the statements within a single version of the code (ferrante1987PDG; le2014patch). However, PDG can only describe the dependency relationships between statements of individual program.
To simultaneously reflect statement dependencies before and after program modification, we first considered the capabilities of the Multi-version Program Dependency Graph ($\delta$-PDG) (parțachi2020flexeme; kim2006program), which is conceptually defined as the disjoint union of all nodes and edges from the PDGs of versions $i$ and $j$. We utilized this structural concept to construct our Commit-Level Merged Graph ($G_{Merged}$). However, $G_{Merged}$ is often massive, presenting a critical challenge for context-aware analysis with LLMs due to input length constraints, leading to inefficient token utilization and potentially diffusing the model’s focus away from the core changes. Therefore, we further explore the specialized extraction of contexts from $G_{Merged}$ to derive representations optimized for LLM consumption, namely the Explicit Context ($G_{Explicit}$) and the Implicit Context ($G_{Implicit}$).
We first construct the Commit-Level Merged Graph ($G_{Merged}$), which serves as the consolidated foundation for our context extraction. As illustrated in Algorithm 1, the construction process begins by taking the PDGs before ($\mathcal{G}_{B}$) and after ($\mathcal{G}_{A}$) the code change for every affected file, along with the commit’s line-level difference ($\mathcal{D}$). Using $\mathcal{D}$, we mark nodes in $\mathcal{G}_{B}$ corresponding to deleted lines as Deleted, and nodes in $\mathcal{G}_{A}$ corresponding to added lines as Added.
Following this initial marking, all file-level marked PDGs are aligned and merged into the single $G_{Merged}$ structure, where common unchanged nodes are consolidated using a fuzzy matching strategy (Lines 1-8). This construction relies on a robust fuzzy node matching strategy to identify semantically equivalent nodes between the two versions. Specifically, a node from $\mathcal{G}_{B}$ is matched with one from $\mathcal{G}_{A}$ if their node type, their structural context similarity, and their label content similarity all exceed defined fuzziness cutoffs ($\tau$), which is set to 1.0 following the practice of Flexeme (parțachi2020flexeme). Nodes successfully matched are consolidated as Unchanged Nodes. The resulting $G_{Merged}$ is constructed by consolidating the matched nodes (Unchanged Nodes) and their edges, retaining the unmatched nodes from $\mathcal{G}_{B}$ (Deleted nodes) and $\mathcal{G}_{A}$ (Added nodes), along with all their respective edges. This approach yields $G_{Merged}$, which represents the disjoint union of nodes and edges across both versions, with common elements consolidated, thereby effectively capturing the complete dependency scope of the committed changes.
Algorithm 1 Explicit and Implicit Contexts Generation
1: $\mathcal{D}$: Commit diff list (lines added/deleted); $\mathcal{G}_{B},\mathcal{G}_{A}$: PDGs of modified files (Before, After); $\tau$: Node matching fuzziness cutoffs.
2: $G_{Explicit}$: Explicit Context.
3: $G_{Implicit}$: Implicit Context.
[…]
5:for each modified file $F$ do
6:  $\mathcal{D}_{F}\leftarrow\text{DIFF\_FOR\_FILE}(\mathcal{D},F)$
7:  $G_{B,F}^{\prime}\leftarrow\text{MARK\_PDG}(\mathcal{G}_{B}[F],\mathcal{D}_{F})$ // Mark deleted nodes according to commit diffs
8:  $G_{A,F}^{\prime}\leftarrow\text{MARK\_PDG}(\mathcal{G}_{A}[F],\mathcal{D}_{F})$ // Mark added nodes according to commit diffs
9:  $G_{\Delta,F}\leftarrow\text{MERGE\_MARKED}(G_{B,F}^{\prime},G_{A,F}^{\prime},\tau)$ // Generate Merged PDG for each file
10:  $\mathcal{L}_{files}\leftarrow\mathcal{L}_{files}\cup\{G_{\Delta,F}\}$
11:end for
[…]
17:for each pair $(S,T)$ in $N_{Changed}\times N_{Changed}$ do // Explicit Context Generation
18:  if $S\neq T$ and $\text{PATH\_EXISTS}(G_{Merged},S,T)$ then
19:   $\text{ADD\_EDGE}(G_{Explicit},S,T,\text{type=`shortcut'})$
20:  end if
21:end for
22:return $G_{Explicit},G_{Implicit}$
For most modification operations, the changed parts account for only a small proportion of the entire program. To reduce unnecessary information, we introduce the concept of explicit contexts. The Explicit Context (i.e., $G_{Explicit}$) is a highly compressed graph designed to capture the logical dependency flow solely among the modified statements, disregarding intervening unchanged code. To construct $G_{Explicit}$, we first initialize a subgraph containing only the changed nodes ($N_{\text{Changed}}$), as shown in Algorithm 1 (Line 10). We then apply dependency compression by adding shortcut edges: for every distinct pair of nodes $(S,T)\in N_{\text{Changed}}$ that have a dependency path in the complete $G_{Merged}$, a direct shortcut edge is added between $S$ and $T$ in $G_{Explicit}$ (Lines 14-18). Although $G_{Explicit}$ loses all surrounding context, it effectively preserves the critical data and control dependencies between the modified statements, offering a minimal, dependency-centric view of the change impact.
Conversely, the Implicit Context (i.e., $G_{Implicit}$ ) is designed to retain neighbor code information without applying compression. $G_{Implicit}$ is generated by extracting the subgraph that includes all changed nodes ($N_{\text{Changed}}$) and their direct one-hop neighbors in $G_{Merged}$ (Lines 11-12). $G_{Implicit}$ strictly preserves only the original PDG edges present among these nodes and contains no shortcut edges. By including the immediate neighbors, $G_{Implicit}$ offers a larger, more comprehensive view of the surrounding code and structural environment than $G_{Explicit}$, enabling LLMs to leverage richer semantic information for accurate analysis.
Figure 3. Code Change Example with Explicit and Implicit Contexts
@@
Figure 3 illustrates a code change scenario involving two statement replacements (at lines 2 and 6). Figure 3 (a) shows the $G_{Merged}$ of this code snippet, where red nodes represent the statements before the change, green nodes represent the modified statements, and gray nodes denote the unchanged statements. Figure 3 (b) presents the explicit contexts ($G_{explicit}$), which are obtained by compressing the $G_{Merged}$ in Figure 3 (a). Specifically, we retain the changed nodes and their direct dependency edges; additionally, if two changed nodes lack a direct dependency but are connected via a dependency path (channel), we add a compressed edge between them. For example, nodes 2 and 6 have no direct dependency but are connected via a path through nodes 4 and 5, resulting in a compressed edge between them. This compressed edge is labeled with the variables involved in the data flow along the original dependency path. Figure 3 (c) shows the implicit contexts ($G_{implicit}$), which retain both the changed nodes and their immediate neighbors (within one-hop dependencies). In this example, node 9 is excluded because it does not have a direct dependency with any of the changed nodes.
@@ 3.2.1. Agent Building
(a) Worker Agent
(b) Reviewer Agent
Figure 4. Agents in Multi-Agent Collaboration
@@
ColaUntangle builds three types of agents. The first two are worker agents, each specializing in untangling commits based on either explicit or implicit dependencies between code changes, as illustrated in Figure 4(a). The third is a reviewer agent, which specializes in reviewing and synthesizing the results from the worker agents, as shown in Figure 4(b).
@@
Worker Agent. Based on the explicit and implicit dependency rules summarized in Section 2.2, we define two types of worker agents, as illustrated in Figure 4(a). The worker agents specialize in untangling commits by analyzing explicit and implicit dependencies between code changes, respectively. We refer to the explicit worker agent as EA and the implicit worker agent as IA. We define the roles of these agents using the following role definition prompt: “You are an expert in untangling tangled commits, specializing in analyzing explicit (or implicit) dependencies between code changes and comprehending their underlying semantic relationships”. We also provide each agent with the corresponding explicit or implicit dependency rules (concluded in Section 2.2) to guide their analysis. Each worker agent has two main functions. The first function is “untangle”. EA untangles tangled commits based on the inputs of commit diffs ($\mathcal{D}$) and explicit contexts ($G_{Explicit}$), while IA performs untangling based on commit diffs ($\mathcal{D}$) and implicit contexts ($G_{Implicit}$). Both agents are required to generate explanations for their decisions to facilitate understanding of their underlying criteria. The second function is “validate”. In this function, each worker agent compares the synthesized result generated by the reviewer agent (described in the next section) with its own untangling result and provides its opinion on the synthesized result. This includes indicating agreement or disagreement and, if disagreeing, providing a revised untangling result along with an explanation.
@@
Reviewer Agent. The untangling results produced by the two worker agents may conflict due to their different expertise and the presence of diverse explicit and implicit dependencies within code changes. To resolve these conflicts, we define a reviewer agent, referred to as RA, which specializes in reviewing and synthesizing the untangling results from the worker agents, as illustrated in Figure 4(b). We define the reviewer agent’s role using the following role definition prompt: “You are an expert reviewer specializing in reviewing and synthesizing untangling results from worker agents. You have deep expertise in analyzing both explicit and implicit dependencies between code changes”. To support its role, we provide RA with both explicit and implicit dependency rules to guide its synthesis process. RA has three main functions: synthesize, revise and stop. In the synthesize function, RA synthesizes the results from EA and IA based on the structured code change information ($\mathcal{D}$, $G_{Explicit}$, and $G_{Implicit}$). In the revise function, RA revises its synthesized result based on the new opinions generated by the two worker agents during validation. Finally, in the stop function, when the collaborative consultation process reaches its maximum iteration limit, RA produces the final untangling result. For each function, RA is required to provide clear explanations for its decisions.
3.2.2. Untangling Workflow
In the first phase, generation of initial reviews, EA and IA are instructed to generate their preliminary untangling results $r_{EA}$ and $r_{IA}$ using their Untangle function, based on the commit diffs ($\mathcal{D}$) and the explicit ($G_{Explicit}$) or implicit ($G_{Implicit}$) contexts:
$r_{EA}=EA.Untangle(Prompt(\mathcal{D},G_{Explicit}))$
$r_{IA}=IA.Untangle(Prompt(\mathcal{D},G_{Implicit}))$
$r_{RA}=RA.Synthesize(Prompt(r_{EA},r_{IA},\mathcal{D},G_{Explicit},G_{Implicit}))$
@@ 3.2.2. Untangling Workflow
1:Experts EAEA, IAIA, RARA; initial synthesized result $r_{RA}$; maximum attempts $t$; prompt template PromptPrompt
@@
[…]
4:while $consensus=False$ and $round<t$ do // Iterative consultation
5:  $round\leftarrow round+1$
6:  $consensus\leftarrow True$
7:  $EA\_opinion\leftarrow EA.validate(Prompt(r_{RA}))$
8:  $IA\_opinion\leftarrow IA.validate(Prompt(r_{RA}))$
9:  if EA_opinion.agree=FalseEA\_opinion.agree=False or IA_opinion.agree=FalseIA\_opinion.agree=False then
10:   $consensus\leftarrow False$
11:   if $round=t$ then
12:     $r_{RA}\leftarrow RA.stop(EA\_opinion,IA\_opinion)$
13:   else
14:     $r_{RA}\leftarrow RA.revise(EA\_opinion,IA\_opinion)$
15:   end if
16:  end if
17:end while
[…]
19:return $r_{f}$
@@
$r_{RA}^{j}=RA.revise(Prompt(EA\_opinion^{j},IA\_opinion^{j}))$
@@
$r_{RA}^{j}=RA.stop(Prompt(EA\_opinion^{j},IA\_opinion^{j}))$
@@ 4.1. Datasets
We use two datasets that have been widely used in previous work: 1) A C# dataset used in (chen2022attributedgraph; herzig2016impact; parțachi2020flexeme; li2022utango; fan2024HD-GNN), which contains 1,612 tangled commits (each of which has $\leq$ 3 concerns) across 9 GitHub projects. 2) A Java dataset used in (li2022utango; fan2024HD-GNN). The Java dataset contains over 14k tangled commits in 10 GitHub repositories, where the number of concerns in a commit ranges from 2-32. Both datasets were constructed by first selecting atomic commits based on criteria such as temporal proximity, namespace similarity, and file coupling, and then artificially tangling them via git cherry-picking to simulate realistic tangled commit scenarios.
@@ 4.2. Baselines
  • Barnett et al. (barnett2015Heuristic-rule): A heuristic rule-based approach considering def-use, use-use, and same-enclosing-method relations among code changes.
  • Herzig et al. (herzig2016impact): A heuristic rule-based approach combining various Confidence Voters and build triangle partition matrix to untangle the commits.
  • Flexeme (parțachi2020flexeme): A graph clustering-based approach builds a $\delta$-NFG (proposed in Flexeme) of commits and then applies agglomerative clustering on this $\delta$-NFG to untangle the commits.
  • $\delta$-PDG+CV (parțachi2020flexeme): A variant of Flexeme by applying Herzig et al.’s confidence voters directly to $\delta$-PDG (proposed in Flexeme).
  • UTango (li2022utango): A Graph clustering-based approach builds a $\delta$-PDG from commits and then applies GNN and agglomerative clustering on this $\delta$-PDG to untangle the commits.
  • HD-GNN (fan2024HD-GNN): A Graph clustering-based approach builds an entity reference graph from commits and then applies GNN to aggregate the embeddings of nodes and edges in the graph and classify the relation between the code changes.
The baselines used for Java dataset are: Barnett et al. (barnett2015Heuristic-rule), Herzig et al. (herzig2016impact), UTango (li2022utango), and HD-GNN (fan2024HD-GNN) mentioned above, as all these approaches support analysis of Java program files. SmartCommit (shen2021smartcommit) is a graph clustering-based interactive method that builds a Diff Hunk Graph and applies a graph partitioning algorithm to untangle commits. We also include SmartCommit and its baselines, Base-1 (a rule-based approach that groups all changes into a single concern) and Base-2 (which groups changes by file), in our comparison.
LLM-based Baselines. We incorporate baselines utilizing single LLM for both C# and Java datasets. We employ two different prompt strategies: zero-shot and zero-shot CoT. Zero-shot refers to prompting LLM to perform a task without providing examples, zero-shot CoT directly incorporates the prompt Let’s think step by step after a question to facilitate inference (kojima2022zeroshot).